# references.bib

### v2

@@ 5 Discussion
In addition, OCP exhibits several unique characteristics that must be accounted for to extract the most performance out of any given model. One such property is given by periodic boundary conditions (common for crystal systems), requiring a dedicated pre-processing step. Another characteristic is the presence of 3 types of atoms: sub-surface slab atoms, surface slab atoms, and adsorbate atoms, which must be distinguished by the model. Finally, the task chosen in the benchmarking Graphormer paper is not only relaxed energy prediction, but also relaxed structure prediction, entailing the prediction of new coordinates for all atoms. This again leverages additional data in the dataset and can be considered a task with synergistic positive effects for relaxed energy prediction.
@@
On top of this, the 3D implementation of Graphormer is unavailable on huggingface (the version used throughout the paper). We chose to perform experiments using a 10K training set that is provided as part of OCP, and the same validation set of size 25K. The entire OCP dataset consists of around 500K dense catalytic structures and training both Graphormer and ESA/NSA on this task would entail a computational effort larger than for any other evaluated dataset. To complicate things further, each ‘graph’ is dense, leading to a significant GPU memory burden, such that only small batch sizes are possible even for high-end GPUs.
@@
Moreover, with higher batch sizes we have hit a limit of PyTorch: the nonzero() tensor method that is used as part of the mask computation is not defined for tensors with more elements than the 32-bit integer limit. While this operation can be chunked in smaller tensors, this induces a significant slowdown while training. Overall, this software limitation highlights the fact that current libraries are not optimised for masked attention. We present other software limitations in SI 5, as well as possible solutions, and we believe that ESA can be significantly optimised with careful software (and even hardware) design.
@@
For the 10K train + 25K validation task, we have adapted our method to use the 3D pre-processing described above, and modified Graphormer to perform only relaxed energy prediction (i.e., without relaxed structure prediction). We used a batch size of 16 for both models, and roughly equivalent settings between NSA and Graphormer, where possible, including 4 layers, 16 attention heads, an embedding/hidden size of 256, and the same learning rate (1e-4). Both methods used mixed precision training. Here, we used the Graphormer 3D implementation from the official repository. We also note that structural information (in the form of edge_index tensors in PyTorch Geometric) is provided in the OCP dataset. They are derived in a similar way to PyTorch Geometric’s radius_graph() function, which connects points based on a distance cutoff. We can use this information for ESA/NSA.
Supplementary Information SI 10 Experimental platform
@@
Representative versions of the software used as part of this paper include Python 3.11, PyTorch version 2.5.1 with CUDA 12.1, PyTorch Geometric 2.5.3 and 2.6.0, PyTorch Lightning 2.4.0, huggingface transformers version 4.35.2, and xFormers version 0.0.27. It is worth noting that attention masking and efficient implementations of attention are early features that are advancing quickly. This means that their behaviour might change unexpectedly and there might be bugs. For example, PyTorch 2.1.1 recently fixed a bug that concerned non-contiguous custom attention masks in the scaled_dot_product_attention function.
@@
Supplementary Information SI 11 Dataset statistics
We present a summary of all the used datasets, along with their size and the maximum number of nodes and edges encountered in a graph in the dataset (Table 11). The last two are important as they determine the shape of the mask and of the inputs for the attention blocks. Technically, we require that the maximum number of nodes/edges is determined per batch and the tensors to be padded accordingly. This per-batch maximum is lower than the dataset maximum for most batches. However, certain operations such as layer normalisation, if performed over the last two dimensions, require a constant value. To enable this, we use the dataset maximum.

### v3

@@ Acknowledgements
In addition, OCP exhibits several unique characteristics that must be taken into account to extract the most performance from any given model. One such property is given by periodic boundary conditions (common for crystal systems), requiring a dedicated pre-processing step. Another characteristic is the presence of 3 types of atoms: sub-surface slab atoms, surface slab atoms, and adsorbate atoms, which must be distinguished by the model. Finally, the task chosen in the benchmarking Graphormer paper is not only relaxed energy prediction but also relaxed structure prediction, entailing the prediction of new coordinates for all atoms. This again leverages additional data in the dataset and can be considered a task with synergistic positive effects for relaxed energy prediction.
@@
On top of this, the 3D implementation of Graphormer is not available on huggingface (the version used throughout the paper). We chose to perform experiments using a 10K training set that is provided as part of OCP, and the same validation set of size 25K. The entire OCP dataset consists of around 500K dense catalytic structures, and training both Graphormer and ESA/NSA on this task would entail a computational effort larger than for any other evaluated dataset. To complicate things further, each structure is dense, leading to a significant GPU memory burden such that only small batch sizes are possible even for high-end GPUs.
@@
Moreover, with higher batch sizes we have hit a limit of PyTorch: the nonzero() tensor method that is used as part of the mask computation is not defined for tensors with more elements than the 32-bit integer limit. Although this operation can be chunked using smaller tensors, this induces a significant slowdown during training. Overall, this software limitation highlights the fact that current libraries are not optimised for masked attention. We present other software limitations in SI 7, as well as possible solutions, and we believe that ESA can be significantly optimised with careful software (and even hardware) design.
@@
For the 10K train + 25K validation task, we have adapted our method to use the 3D pre-processing described above, and modified Graphormer to perform only relaxed energy prediction (i.e. without relaxed structure prediction). We used a batch size of 16 for both models, and roughly equivalent settings between NSA and Graphormer, where possible, including 4 layers, 16 attention heads, an embedding/hidden size of 256, and the same learning rate (1e-4). Both methods used mixed precision training. Here, we used the Graphormer 3D implementation from the official repository. We also note that structural information (in the form of edge_index tensors in PyTorch Geometric) is provided in the OCP dataset. They are derived in a similar way to PyTorch Geometric’s radius_graph() function, which connects points based on a distance cutoff. We can use this information for ESA/NSA.
Supplementary Information SI 13 Experimental platform
@@
Representative versions of the software used as part of this paper include Python 3.11, PyTorch version 2.5.1 with CUDA 12.1, PyTorch Geometric 2.5.3 and 2.6.0, PyTorch Lightning 2.4.0, huggingface transformers version 4.35.2, and xFormers version 0.0.27. It is worth noting that attention masking and efficient implementations of attention are early features that are advancing rapidly. This means that their behaviour might change unexpectedly and there might be bugs. For example, PyTorch 2.1.1 recently fixed a bug that concerned non-contiguous custom attention masks in the scaled_dot_product_attention function.
@@
Supplementary Information SI 14 Dataset statistics
We present a summary of all the used datasets, together with their size and the maximum number of nodes and edges encountered in a graph in the dataset (Table 15). The last two are important, as they determine the shape of the mask and of the inputs for the attention blocks. Technically, we require the maximum number of nodes/edges to be determined per batch, and the tensors to be padded accordingly. This per-batch maximum is lower than the dataset maximum for most batches. However, certain operations such as layer normalisation, if performed over the last two dimensions, require a constant value. To enable this, we use the dataset maximum.
Supplementary Information SI 15 Sourcing and licensing
@@
Supplementary Information SI 16 Explainability
We computed ground truth molecular orbitals using Psi4 and the same methodology as in the original QM9 paper (B3LYP functional with the 6-31G(d) basis set) [Ramakrishnan2014]. The ground truth is in this case an approximation given by visualisation parameters such as the isosurface value. A lower isosurface value includes regions of lower electron density and thus a more extended/diffuse surface, while higher values would highlight only the regions with the highest density. Here, we use the default parameters in Avogadro 2.
It is worth noting that our models are trained on 2D graph representations without 3D information, and only to predict the scalar HOMO energy. As such, it would be unreasonable to expect high-quality 3D molecular orbital reconstructions. Furthermore, attention scores should be regarded as a proxy and potential explainability tool, but it is difficult if not impossible to establish a strong and reliable relationship between the actual physical coefficients and the model weights.
Nonetheless, we visualise the highest attention weights and the HOMO orbitals in Figure 3 for several molecules and notice that the bonds and underlying atoms highlighted by attention are contained within the highlighted 3D regions. The visualisation includes molecules with varied structures and sizes. It is interesting to note that in the case of gdb_39511 (Figure 3C), the molecule is symmetric as far as the 2D model is concerned, and two pairs are highlighted with equal scores. The actual 3D geometry has a different structure, but the highlighted bond corresponds to the visualised orbital.
Supplementary Figure 3: A – D. Top attention weights illustrated through 2D molecular graphs and HOMO visualisations of the corresponding actual 3D geometry from the qm9 dataset. The labels for the four visualised molecules are gdb_117145 (A), gdb_34 (B), gdb_39511 (C), gdb_31595 (D).
Supplementary Information SI 17 Line graphs for graphs indistinguishable by 1-WL
Figure 4 shows an example of two graphs, $G_{1}$ and $G_{2}$, which are indistinguishable by the 1-WL test, for example as provided by the weisfeiler_lehman_graph_hash function from the networkx library. However, the resulting line graphs ($L_{1}$ and $L_{2}$) are now distinguishable by the 1-WL test.
Supplementary Figure 4: Two graphs that are indistinguishable by the 1-WL test ($G_{1}$ and $G_{2}$), while their line graphs counterparts ($L_{1}$ and $L_{2}$) are distinguishable.
[…]
[resetnumbers=true,filter=appendixOnlyFilter,title=Supplementary References]